# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ Abstract
Recent generative reasoning breakthroughs have transformed how large language models (LLMs) tackle complex problems by dynamically retrieving and refining information while generating coherent, multi-step thought processes. Techniques such as inference-time scaling, reinforcement learning, supervised fine-tuning, and distillation have been successfully applied to models like DeepSeek-R1, OpenAI’s o1 & o3, GPT-4o, Qwen-32B, and various Llama variants, resulting in enhanced reasoning capabilities. In this paper, we provide a comprehensive analysis of the top 27 LLM models released between 2023 and 2025 (including models such as Mistral AI Small 3 24B, DeepSeek-R1, Search-o1, QwQ-32B, and phi-4). Then, we present an extensive overview of training methodologies that spans general training approaches, mixture-of-experts (MoE) and architectural innovations, retrieval-augmented generation (RAG), chain-of-thought and self-improvement techniques, as well as test-time compute scaling, distillation, and reinforcement learning (RL) methods. Finally, we discuss the key challenges in advancing LLM capabilities, including improving multi-step reasoning without human supervision, overcoming limitations in chained tasks, balancing structured prompts with flexibility, and enhancing long-context retrieval and external tool integration.
@@ I Introduction
The recent emergence of large reasoning models (LRMs) such as OpenAI’s o1 [1], QwenQwQ [2], and DeepSeek-R1 [3] marks a significant milestone in AI research. These models leverage large-scale reinforcement learning to achieve impressive long-sequence, stepwise reasoning capabilities. Their ability to break down complex problems into extended chains of thought has inspired numerous foundational efforts to replicate and broaden these reasoning patterns across various model architectures [4]. A key characteristic of these o1-like reasoning patterns is their promoting a slower, more deliberate thinking process. By decomposing intricate problems into stepwise reasoning chains, LRMs can enhance logical coherence and interpretability. However, this extended reasoning process also introduces challenges such as overthinking and the risk of propagating errors through knowledge gaps, which may disrupt the reasoning chain [5].
@@
Mathematical reasoning remains a formidable challenge in AI, with applications ranging from automated theorem proving to scientific discovery [6]. Recent advances, such as those seen in GPT-4, have improved performance in arithmetic and geometric tasks, yet complex problems, especially at Olympiad-level benchmarks, continue to test these models [7, 8]. Researchers have explored breaking down solutions into step-by-step reasoning paths—using methods like Chain-of-Thought [9], Self-Refine [10], and Reflexion—to enhance problem-solving [11], although these approaches sometimes face challenges related to feedback inefficiencies and local optima. Simultaneously, effective evaluation methods, including outcome and process reward models [12], struggle with inconsistent labeling and varied scoring standards. Innovations integrating techniques like Monte Carlo Tree Search with reinforcement learning-based evaluations offer promising avenues to improve the efficiency and reliability of mathematical reasoning in AI [13].
@@
Large language models (LLMs) have achieved impressive performance on complex reasoning tasks through prompting methods that induce chain-of-thought (CoT) reasoning, but recent findings reveal that by simply modifying the decoding procedure—specifically by inspecting alternative top-k tokens—LLMs can naturally elicit CoT reasoning without additional prompts or tuning [14, 15, 16]. Therefore, text often conceals meaning in its implications, and while language models improve with reasoning, traditional methods like curated QA datasets (as used in Self-Taught Reasoner, or STaR) [17] limit their scalability and generality. Quiet-STaR [18] extends this approach by training LMs on a vast internet corpus using REINFORCE-based rewards, enabling them to learn diverse, implicit reasoning skills and ”think before speaking” without relying solely on manually curated data.
@@
Traditional LLMs lack inherent ”Thinking then Responding” abilities because their training data often omits explicit thinking process annotations. Recent efforts address this gap by generating synthetic thinking data—using techniques like Monte Carlo Tree Search [19], GPT-4o demonstrations [20], and model distillation [21]—or by employing reinforcement learning [22] to enable autonomous reasoning path learning. Despite strong benchmark performances, LLMs still face challenges in advanced reasoning tasks like mathematical problem-solving and code generation [23]. Techniques such as chain-of-thought prompting [9], fine-tuning with CoT examples [16], and RLHF [24] have enhanced reasoning capabilities, yet complex multi-step problems remain problematic. While decoding strategies like self-consistency and verifiers help, reward models—especially Process Reward Models offering step-by-step feedback—prove more effective but are hindered by costly, manual supervision [25]. To address this, OmegaPRM [26] introduces a divide-and-conquer Monte Carlo Tree Search algorithm that automates the collection of over 1.5 million high-quality process annotations, significantly boosting performance on benchmarks such as MATH500 and GSM8K.
@@
This paper investigates strategies for enhancing the reasoning capabilities of large language models by integrating methods such as inference-time scaling, pure reinforcement learning, and combined supervised fine-tuning with reinforcement learning, as well as distillation. We examine how models like DeepSeek-R1, OpenAI’s o1, GPT-4o, Qwen-32B, and Llama variants are refined to generate transparent, multi-step reasoning processes that improve performance on complex tasks ranging from advanced mathematics to coding challenges. The main contributions of this study are:
@@
Fig. 1 illustrates this survey’s structure. Section II presents the related works. Section III analyzes the top 27 LLM models released from 2023 to 2025. Section IV overviews training methodologies for LLMs, categorizing them into general approaches, MoE and architectural innovations, RAG, chain-of-thought and self-improvement, test-time scaling and distillation, and RL methods. Section V discusses key challenges in advancing LLM capabilities. Finally, Section VI concludes the paper.
@@
Yin et al. [27] 2024 Multimodal LLMs Architecture details, training strategies, M-ICL, M-CoT, LAVR Overview of MLLMs with novel capabilities (e.g., narrative generation, OCR-free math reasoning) Partial Yes Partial
Plaat et al. [28] 2024 LLM Reasoning In-context learning, chain-of-thought prompting Develops a taxonomy of reasoning strategies with insights into multi-step inference challenges No Partial Yes
Wang et al. [29] 2024 RL Integration with LLMs RL-enhanced architectures, reward modeling (RLHF, RLAIF, DPO) Reviews integration challenges and reward modeling strategies for LLMs No Yes Yes
Wang et al. [30] 2024 RL for Code Generation RL in compiler optimization, resource allocation, code generation frameworks Optimizes code generation and system efficiency via RL-based approaches No Yes Partial
Cao et al. [31] 2024 LLMs in RL Roles: information processor, reward designer, decision-maker, generator Enhances RL via LLM integration; addresses multi-task learning, sample efficiency, and high-level task planning No Partial Partial
Wang et al. [32] 2024 LLM Alignment Techniques RLHF, RLAIF, PPO, DPO Categorizes alignment methods to improve reliability and safety of LLM outputs No Yes Partial
Kumar et al. [33] 2025 Post-training for LLMs Fine-tuning, reinforcement learning, testing-time scaling Enhances reasoning, factual accuracy, and adaptability; tackles issues like catastrophic forgetting No Yes Yes
Tie et al. [34] 2025 Post-training Methodologies Fine-tuning, alignment, reasoning, efficiency, integration/adaptation Provides a structured taxonomy of post-training evolution No Yes Yes
Yigit et al. [35] 2025 Cybersecurity & LLMs Evaluation benchmarks for cybersecurity, agentic AI for proactive defense Secures critical infrastructures by addressing trust, privacy, and resilience No No Yes
Chen et al. [36] 2025 Long Chain-of-Thought Reasoning Novel taxonomy of reasoning paradigms, deep reasoning, reflective processes Differentiates Long CoT from traditional approaches and discusses integration with multimodal reasoning No Partial Yes
Yehudai et al. [37] 2025 Evaluation of LLM-based Agents Systematic analysis of evaluation benchmarks and frameworks across agent capabilities, application-specific benchmarks, generalist agents, and evaluation methodologies First comprehensive survey on evaluation methodologies for LLM-based agents; highlights trends toward more realistic, challenging evaluations and identifies gaps in cost-efficiency, safety, and scalability No No Yes
Ferrag et al. [38] 2025 Cybersecurity and Generative AI LLM applications across hardware security, intrusion detection, software engineering, design verification, cyber threat intelligence, malware and phishing detection Provides a comprehensive review on integrating generative AI and LLMs into cybersecurity frameworks Partial Partial No
@@ II Related Works
Recent LLM surveys have increasingly sought to encapsulate the rapidly evolving landscape of methodologies and applications. For example, several studies from 2024 have focused on specialized domains such as multimodal LLMs, reasoning strategies, and reinforcement learning integrations, while surveys from 2025 have extended these analyses to include cybersecurity applications and long chain-of-thought reasoning, as presented in Tab. I. These surveys systematically address various aspects, including architectural innovations, training strategies (e.g., fine-tuning, reinforcement learning, and in-context learning), and evaluation benchmarks.
@@ II-A Multimodal Applications and Cybersecurity
Yin et al. [27] provides an in-depth overview of the rapidly evolving field of multimodal large language models (MLLMs), exemplified by cutting-edge systems like GPT-4V. The authors lay a solid foundation by describing the core formulation of MLLMs, including their architectures, training strategies, data usage, and evaluation metrics. They emphasize the unique capabilities of these models—such as generating narratives from images and performing OCR-free mathematical reasoning—which are seldom observed in traditional multimodal approaches and hint at the potential trajectory toward artificial general intelligence. The review further explores avenues for enhancing MLLMs by extending support to additional granularities, modalities, languages, and application scenarios. It also tackles critical issues like multimodal hallucination and introduces innovative techniques, including Multimodal In-Context Learning (M-ICL), Multimodal Chain-of-Thought (M-CoT), and LLM-Aided Visual Reasoning (LAVR), to improve the models’ reasoning capabilities.
The review by Yigit et al. [35] provides a comprehensive examination of how generative AI and LLMs can be harnessed to bolster the cybersecurity of critical national infrastructures such as energy grids, water systems, transportation networks, and communication frameworks. It begins by assessing the reliability of these infrastructures using established evaluation benchmarks tailored for cybersecurity applications of LLMs, and it then delves into fundamental issues like trust, privacy, resilience, and securability. Notably, the paper highlights the emerging role of agentic AI in developing proactive defense mechanisms, thereby offering innovative pathways for detecting and mitigating cyber threats. By synthesizing current challenges with advanced AI methodologies, the authors propose a strategic roadmap that reinforces existing protection measures and sets the stage for future research in safeguarding critical infrastructures.
The survey by Ferrag et al. [38] offers a comprehensive review of integrating generative AI and large language models (LLMs) within cybersecurity frameworks. It systematically examines how LLMs, including state-of-the-art models such as GPT-4, GPT-3.5, Mixtral-8x7B, BERT, Falcon2, and LLaMA, can be applied across various cybersecurity domains—from hardware design security and intrusion detection to software engineering and malware as well as phishing detection. The study not only highlights the evolution and current state of LLMs but also critically assesses their vulnerabilities, including prompt injection, insecure output handling, data poisoning, DDoS attacks, and adversarial instructions. Furthermore, it evaluates the performance of 42 distinct LLM models in terms of cybersecurity knowledge and hardware security, offering valuable insights into their strengths and limitations.
@@ II-B Reinforcement Learning
Cao et al. [31] The paper comprehensively reviews how LLMs can enhance reinforcement learning (RL) by integrating extensive pre-trained knowledge into the learning process. It systematically categorizes the contributions of LLMs into four distinct roles—information processor, reward designer, decision-maker, and generator—each designed to address specific challenges in RL, such as multi-task learning, sample efficiency, and high-level task planning. By framing its analysis within the classical agent-environment interaction paradigm, the survey not only highlights the methodologies underpinning each role but also contrasts these innovations with conventional RL approaches. Furthermore, the paper presents a comparative analysis that elucidates potential applications in complex domains like robotics, autonomous driving, and energy systems, and it outlines future research directions aimed at accelerating progress in the field.
Wang et al. [29] provides a comprehensive overview of how reinforcement learning (RL) can be integrated with LLMs to significantly enhance their performance, as illustrated by models like DeepSeek-R1. The authors detail the complexities involved in this integration, including developing sophisticated algorithms, reward modeling strategies, and optimization techniques critical for effective RL implementations. The survey systematically reviews the fundamentals of RL, introduces leading RL-enhanced LLM architectures, and critically examines two primary reward model-based methods: reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF), as well as direct preference optimization (DPO), which leverages human preference data directly. By identifying the current challenges and gaps in existing methodologies, the paper lays out a strategic agenda for future research, aiming to foster deeper understanding and further advancements in the field.
Wang et al. [30] examine the integration of reinforcement learning (RL) techniques with LLMs for code generation and optimization. It highlights how RL has become pivotal in enhancing compiler optimization, where sophisticated algorithms are employed to improve efficiency and resource utilization, particularly in tasks like register allocation and overall system optimization. In addition, the paper discusses RL’s role in advancing resource allocation strategies and strengthening the development of robust code generation frameworks and tools. By providing a detailed review of these applications, the survey serves as a comprehensive resource for researchers and practitioners seeking to leverage RL to refine and innovate code generation methodologies, ultimately addressing both current challenges and potential future directions in the field.

### v2

@@ Abstract
Recent breakthroughs in generative reasoning have fundamentally reshaped how large language models (LLMs) address complex tasks, enabling them to dynamically retrieve, refine, and organize information into coherent, multi-step reasoning chains. Techniques such as inference-time scaling, reinforcement learning, supervised fine-tuning, and distillation have been effectively applied to state-of-the-art models, including DeepSeek-R1, OpenAI’s o1 and o3, GPT-4o, Qwen-32B, and various Llama variants, significantly enhancing their reasoning capabilities. In this paper, we present a comprehensive review of the top 27 LLMs released between 2023 and 2025, such as Mistral AI Small 3 24B, DeepSeek-R1, Search-o1, QwQ-32B, and Phi-4, and analyze their core innovations and performance improvements.
We also provide a detailed overview of recent advancements in multilingual large language models (MLLMs), emphasizing methods that improve cross-lingual reasoning and address the limitations of English-centric training. In parallel, we present a comprehensive review of progress in State Space Model (SSM)-based architectures, including models like Mamba, which demonstrate improved efficiency for long-context processing compared to Transformer-based approaches. Our analysis covers training strategies such as general optimization techniques, mixture-of-experts (MoE) configurations, retrieval-augmented generation (RAG), chain-of-thought prompting, self-improvement methods, and test-time compute scaling and distillation frameworks.
Finally, we identify key challenges for future research, including enabling multi-step reasoning without human supervision, improving robustness in chained task execution, balancing structured prompting with generative flexibility, and enhancing the integration of long-context retrieval and external tools.
@@
Figure 1: Standard LLMs process an input prompt and immediately generate a response. Reasoning-enabled LLMs, by contrast, interpose a chain of internal “thought” steps. Each step builds on the last to work through the problem methodically before producing the final answer.
@@ I Introduction
The recent emergence of large reasoning models (LRMs) such as OpenAI’s o1 [jaech2024openai], QwenQwQ [bai2025qwen2], and DeepSeek-R1 [guo2025deepseek] marks a significant milestone in AI research [saoud2024artificial]. These models leverage large-scale reinforcement learning to achieve impressive long-sequence, stepwise reasoning capabilities. Their ability to break down complex problems into extended chains of thought has inspired numerous foundational efforts to replicate and broaden these reasoning patterns across various model architectures [zhang2025lessons]. A key characteristic of these o1-like reasoning patterns is their promoting a slower, more deliberate thinking process. By decomposing intricate problems into stepwise reasoning chains, LRMs can enhance logical coherence and interpretability. However, this extended reasoning process also introduces challenges such as overthinking and the risk of propagating errors through knowledge gaps, which may disrupt the reasoning chain [li2025search, muhammad2024pre, mohammadi2023ai, ajabani2025enhancing].
@@
Mathematical reasoning remains a formidable challenge in AI, with applications ranging from automated theorem proving to scientific discovery [zhang2024llama, ahmad2025liu, shahid2025blockchain]. Recent advances, such as those seen in GPT-4, have improved performance in arithmetic and geometric tasks, yet complex problems, especially at Olympiad-level benchmarks, continue to test these models [aops_aime, he2024olympiadbench]. Researchers have explored breaking down solutions into step-by-step reasoning paths using methods like Chain-of-Thought [wei2022chain], Self-Refine [madaan2023self], and Reflexion to enhance problem-solving [shinn2023reflexion], although these approaches sometimes face challenges related to feedback inefficiencies and local optima. Simultaneously, effective evaluation methods, including outcome and process reward models [uesato2022solving, liu2025efficient], struggle with inconsistent labeling and varied scoring standards. Innovations integrating techniques like Monte Carlo Tree Search with reinforcement learning-based evaluations offer promising avenues to improve the efficiency and reliability of mathematical reasoning in AI [christiano2017deep, akhund2024comprehensive].
@@
Large language models (LLMs) have achieved impressive performance on complex reasoning tasks through prompting methods that induce chain-of-thought (CoT) reasoning. Still, recent findings reveal that by simply modifying the decoding procedure, specifically by inspecting alternative top-k tokens, LLMs can naturally elicit CoT reasoning without additional prompts or tuning [wang2024chain, chung2024scaling, cobbe2021training, kousar2024mlhs]. Therefore, text often conceals meaning in its implications, and while language models (LMs) improve with reasoning, traditional methods like curated QA datasets (as used in Self-Taught Reasoner, or STaR) [zelikman2022star, yang2024optimizing] limit their scalability and generality. Quiet-STaR [zelikman2024quiet, shaikh2024new, shaikh2014towards] extends this approach by training LMs on a vast internet corpus using REINFORCE-based rewards, enabling them to learn diverse, implicit reasoning skills and ”think before speaking” without relying solely on manually curated data.
@@
Traditional LLMs lack inherent ”Thinking then Responding” abilities because their training data often omits explicit thinking process annotations. Fig. 1 highlights the key difference between “regular” and “reasoning” LLMs. A user’s question is passed straight into the model in the standard setup, emitting an answer in one step. By contrast, a reasoning-enabled LLM breaks the problem into a series of internal “thought” stages, each building on the previous one before arriving at its final response. This multi-step reasoning process makes the model’s decision path more transparent and helps it tackle complex, multi-stage tasks with greater reliability. Recent efforts address this gap by generating synthetic thinking data using techniques like Monte Carlo Tree Search [guan2025rstar], GPT-4o demonstrations [chen2024huatuogpt], and model distillation [muennighoff2025s1] or by employing reinforcement learning [team2025kimi] to enable autonomous reasoning path learning. Despite strong benchmark performances, LLMs still face challenges in advanced reasoning tasks like mathematical problem-solving and code generation [team2024gemini]. Techniques such as chain-of-thought prompting [wei2022chain], fine-tuning with CoT examples [cobbe2021training], and RLHF [ouyang2022training] have enhanced reasoning capabilities, yet complex multi-step problems remain problematic. While decoding strategies like self-consistency and verifiers help, reward models especially Process Reward Models offering step-by-step feedback prove more effective but are hindered by costly, manual supervision [lightman2023let]. To address this, OmegaPRM [luo2024improve] introduces a divide-and-conquer Monte Carlo Tree Search algorithm that automates the collection of over 1.5 million high-quality process annotations, significantly boosting performance on benchmarks such as MATH500 and GSM8K.
@@
This paper investigates strategies for enhancing the reasoning capabilities of LLMs by integrating methods such as inference-time scaling, pure reinforcement learning, and combined supervised fine-tuning with reinforcement learning, as well as distillation. We examine how models like DeepSeek-R1, OpenAI’s o1, GPT-4o, Qwen-32B, and Llama variants are refined to generate transparent, multi-step reasoning processes that improve performance on complex tasks ranging from advanced mathematics to coding challenges. The main contributions of this study are:
@@
  • We present a comprehensive overview of recent advancements in multilingual large language models (MLLMs).
  • We present a comprehensive review of recent advancements in State Space Model (SSM)-based architectures.
@@
Fig. 2 illustrates this survey’s structure. Section II presents the related works. Section III analyzes the top 27 LLM models released from 2023 to 2025.Section IV presents a comprehensive overview of recent advancements in multilingual large language models.Section V presents a comprehensive overview of recent advancements in State Space Model (SSM)-based architectures. Section VI overviews training methodologies for LLMs, categorizing them into general approaches, MoE and architectural innovations, RAG, chain-of-thought and self-improvement, test-time scaling and distillation, and RL methods. Section VII discusses key challenges in advancing LLM capabilities. Finally, Section VIII concludes the paper.
@@
Yin et al. [Yin_2024] 2024 Multimodal LLMs Architecture details, training strategies, M-ICL, M-CoT, LAVR Overview of MLLMs with novel capabilities (e.g., narrative generation, OCR-free math reasoning) Partial Yes Partial
Plaat et al. [plaat2024reasoning] 2024 LLM Reasoning In-context learning, chain-of-thought prompting Develops a taxonomy of reasoning strategies with insights into multi-step inference challenges No Partial Yes
Wang et al. [wang2024reinforcement] 2024 RL Integration with LLMs RL-enhanced architectures, reward modeling (RLHF, RLAIF, DPO) Reviews integration challenges and reward modeling strategies for LLMs No Yes Yes
Wang et al. [wang2024enhancing] 2024 RL for Code Generation RL in compiler optimization, resource allocation, and code generation frameworks Optimizes code generation and system efficiency via RL-based approaches No Yes Partial
Cao et al. [cao2024survey] 2024 LLMs in RL Roles: information processor, reward designer, decision-maker, generator Enhances RL via LLM integration; addresses multi-task learning, sample efficiency, and high-level task planning No Partial Partial
Wang et al. [wang2024comprehensive] 2024 LLM Alignment Techniques RLHF, RLAIF, PPO, DPO Categorizes alignment methods to improve the reliability and safety of LLM outputs No Yes Partial
Kumar et al. [kumar2025llm] 2025 Post-training for LLMs Fine-tuning, reinforcement learning, testing-time scaling Enhances reasoning, factual accuracy, and adaptability; tackles issues like catastrophic forgetting No Yes Yes
Tie et al. [tie2025survey] 2025 Post-training Methodologies Fine-tuning, alignment, reasoning, efficiency, integration/adaptation Provides a structured taxonomy of post-training evolution No Yes Yes
Yigit et al. [yigit2025generative] 2025 Cybersecurity & LLMs Evaluation benchmarks for cybersecurity, agentic AI for proactive defense Secures critical infrastructures by addressing trust, privacy, and resilience No No Yes
Chen et al. [chen2025towards] 2025 Long Chain-of-Thought Reasoning Novel taxonomy of reasoning paradigms, deep reasoning, reflective processes Differentiates Long CoT from traditional approaches and discusses integration with multimodal reasoning No Partial Yes
Yehudai et al. [yehudai2025surveyevaluationllmbasedagents] 2025 Evaluation of LLM-based Agents Systematic analysis of evaluation benchmarks and frameworks across agent capabilities, application-specific benchmarks, generalist agents, and evaluation methodologies First comprehensive survey on evaluation methodologies for LLM-based agents; highlights trends toward more realistic, challenging evaluations and identifies gaps in cost-efficiency, safety, and scalability No No Yes
Ferrag et al. [ferrag2025generative] 2025 Cybersecurity and Generative AI LLM applications across hardware security, intrusion detection, software engineering, design verification, cyber threat intelligence, malware, and phishing detection Provides a comprehensive review on integrating generative AI and LLMs into cybersecurity frameworks Partial Partial No
@@ II Related Works
Recent LLM surveys have increasingly sought to encapsulate the rapidly evolving landscape of methodologies and applications. For example, several studies from 2024 have focused on specialized domains such as multimodal LLMs, reasoning strategies, and reinforcement learning integrations, while surveys from 2025 have extended these analyses to include cybersecurity applications and long chain-of-thought reasoning, as presented in Table. I. These surveys systematically address various aspects, including architectural innovations, training strategies (e.g., fine-tuning, reinforcement learning, and in-context learning), and evaluation benchmarks.
@@ II-A Multimodal Applications and Cybersecurity
Yin et al. [Yin_2024] provide an in-depth overview of the rapidly evolving field of multimodal large language models (MLLMs), exemplified by cutting-edge systems like GPT-4V. The authors lay a solid foundation by describing the core formulation of MLLMs, including their architectures, training strategies, data usage, and evaluation metrics. They emphasize the unique capabilities of these models such as generating narratives from images and performing OCR-free mathematical reasoning which are seldom observed in traditional multimodal approaches and hint at the potential trajectory toward artificial general intelligence. The review further explores avenues for enhancing MLLMs by extending support to additional granularities, modalities, languages, and application scenarios. It also tackles critical issues like multimodal hallucination and introduces innovative techniques, including Multimodal In-Context Learning (M-ICL), Multimodal Chain-of-Thought (M-CoT), and LLM-Aided Visual Reasoning (LAVR), to improve the models’ reasoning capabilities.
The review by Yigit et al. [yigit2025generative] provides a comprehensive examination of how generative AI and LLMs can be harnessed to bolster the cybersecurity of critical national infrastructures such as energy grids, water systems, transportation networks, and communication frameworks. It begins by assessing the reliability of these infrastructures using established evaluation benchmarks tailored for cybersecurity applications of LLMs, and it then delves into fundamental issues like trust, privacy, resilience, and securability. Notably, the paper highlights the emerging role of agentic AI in developing proactive defense mechanisms, thereby offering innovative pathways for detecting and mitigating cyber threats. By synthesizing current challenges with advanced AI methodologies, the authors propose a strategic roadmap that reinforces existing protection measures and sets the stage for future research in safeguarding critical infrastructures.
The survey by Ferrag et al. [ferrag2025generative] offers a comprehensive review of integrating generative AI and large language models (LLMs) within cybersecurity frameworks. It systematically examines how LLMs, including state-of-the-art models such as GPT-4, GPT-3.5, Mixtral-8x7B, BERT, Falcon2, and LLaMA, can be applied across various cybersecurity domains from hardware design security and intrusion detection to software engineering and malware as well as phishing detection. The study not only highlights the evolution and current state of LLMs but also critically assesses their vulnerabilities, including prompt injection, insecure output handling, data poisoning, DDoS attacks, and adversarial instructions. Furthermore, it evaluates the performance of 42 distinct LLM models regarding cybersecurity knowledge and hardware security, offering valuable insights into their strengths and limitations.
@@ II-B Reinforcement Learning
Cao et al. [cao2024survey] comprehensively reviews how LLMs can enhance reinforcement learning (RL) by integrating extensive pre-trained knowledge into the learning process. It systematically categorizes the contributions of LLMs into four distinct roles: information processor, reward designer, decision-maker, and generator, each designed to address specific challenges in RL, such as multi-task learning, sample efficiency, and high-level task planning. By framing its analysis within the classical agent-environment interaction paradigm, the survey not only highlights the methodologies underpinning each role but also contrasts these innovations with conventional RL approaches. Furthermore, the paper presents a comparative analysis that elucidates potential applications in complex domains like robotics, autonomous driving, and energy systems, and it outlines future research directions aimed at accelerating progress in the field.
Wang et al. [wang2024reinforcement] provides a comprehensive overview of how reinforcement learning (RL) can be integrated with LLMs to significantly enhance their performance, as illustrated by models like DeepSeek-R1. The authors detail the complexities involved in this integration, including developing sophisticated algorithms, reward modeling strategies, and optimization techniques critical for practical RL implementations. The survey systematically reviews the fundamentals of RL, introduces leading RL-enhanced LLM architectures, and critically examines two primary reward model-based methods: reinforcement learning from human feedback (RLHF) and reinforcement learning from AI feedback (RLAIF), as well as direct preference optimization (DPO), which leverages human preference data directly. By identifying the current challenges and gaps in existing methodologies, the paper lays out a strategic agenda for future research, aiming to foster deeper understanding and further advancements in the field.
Wang et al. [wang2024enhancing] examine the integration of reinforcement learning (RL) techniques with LLMs for code generation and optimization. It highlights how RL has become pivotal in enhancing compiler optimization, where sophisticated algorithms are employed to improve efficiency and resource utilization, particularly in tasks like register allocation and overall system optimization. In addition, the paper discusses RL’s role in advancing resource allocation strategies and strengthening the development of robust code generation frameworks and tools. By providing a detailed review of these applications, the survey is a comprehensive resource for researchers and practitioners seeking to leverage RL to refine and innovate code generation methodologies, ultimately addressing current challenges and potential future directions in the field.